Let V=R. Define addition and scalar multiplication by a+b=4a+4b. Show whether addition is both commutative and associative.


1
Expert's answer
2021-07-12T12:24:08-0400

"a+b=4a+4b\\
c+d=4c+4d\\
c+d+a+b=4a+4b+4c+4d\\
c+d+a+b=4(c+d)+4(a+b)\\"

Let

"A=c+d\\
B=a+b\\
\therefore\\
A+B=4A+4B"

The above defines addition by a+b=4a+4b

Also,

Let s be a scalar

Therefore,

"s(a+b)=4sa+4sb"

The above defines scalar multiplication by a+b=4a+4b

"a+b=4a+4b\\
b+a=4b+4a\\
b+a=a+b"

Hence addition is commutative

"a+b=4a+4b\\
(a+b)+c=4(a+b)+4c=4a+4b+4c\\
a+(b+c)=4a+4(b+c)=4a+4b+4c\\
(a+b)+c=a+(b+c)"

Hence addition is associative